Todrick Company is a merchandiser that reported the following information based on 1,000 units sold: Sales Beginning merchandReg 1 Req 2 Req 3 to 5 Req 6 Prepare a contribution format income statement. Todrick Company Contribution Format Income StateReq 1 Req 2 Req 3 to 5 Req 6 Prepare a traditional format income statement. Todrick Company Traditional Income Statement SellComplete in questiuII Dy enery yuul duiswers II Req 1 Req 2 Req 3 to 5 Req 6 3. Calculate the selling price per unit. 4. CalcComplete this question by entering your answers in the tabs below. Req 1 Req 2 Req 3 to 5 Req 6 Which income statement format
Answer #1

1.

Cost of goods sold = Beginning merchandise inventory+Purchases-Ending merchandise inventory

Cost of goods sold = $29,000+290,000-14,500 = $304,500

Todrick Company

Contribution Format Income Statement

Sales $435,000
Variable expenses:
Cost of goods sold $304,500
Variable selling expense 21,750
Variable administrative expense ($348,000-304,500-21,750) 21,750
Total variable expense ($435,000-87,000) 348,000
Contribution margin 87,000
Fixed expenses:
Fixed selling expense ($60,900-17,400) 43,500
Fixed administrative expense 17,400
Total fixed expenses ($87,000-26,100) 60,900
Net operating income $26,100

2.

Todrick Company

Traditional Income Statement

Sales $435,000
Cost of goods sold (304,500)
Gross profit 130,500
Selling and administrative expenses:
Variable selling expense $21,750
Fixed selling expense 43,500
Variable administrative expense 21,750
Fixed administrative expense 17,400 (104,400)
Net operating income $26,100

3. Selling price per unit = $435,000 / 1,000 = $435

4. Variable cost per unit = $348,000 / 1,000 = $348

5. Contribution margin per unit = $87,000 / 1,000 = $87

6. Contribution income statement will be more useful to managers.